In the following summary of data for a payroll period, some amounts have been intentionally omitted. Calculate the amounts omitted below. Enter all your answers as positive amounts.

Earnings:
1. At regular rate $
2. At overtime rate 120,000
3. Total earnings $
Deductions:
4. FICA tax $63,000
5. Income tax withheld 200,000
6. Medical insurance 37,000
7. Union dues
8. Total deductions $315,000
9. Net amount paid $525,000
Accounts increased:
10. Factory Wages $400,000
11. Sales Salaries $
12. Office Salaries $150,000
Answer #1

Answers:

3. Total Earnings = Net amount paid + Total deductions

=525,000+315,000

=840,000

1. Regular = Total earnings - At overtime rate

    =840,000 - 120,000

    =720,000

7. Union dues =Total deductions - (FICA tax+Income tax withheld+Medical insurance)

    =315,000 - (63,000+200,000+37,000)

    =15000

11. Sales Salaries = Total earnings - (Factory Wages+Office Salaries)

= 840,000 - (400,000+150,000)

= 290,000

Earnings:
1. At regular rate 720000
2. At overtime rate 120,000
3. Total earnings 840000
Deductions:
4. FICA tax $63,000
5. Income tax withheld 200,000
6. Medical insurance 37,000
7. Union dues $ 15,000
8. Total deductions $315,000
9. Net amount paid $525,000
Accounts increased:
10. Factory Wages $400,000
11. Sales Salaries $290,000
12. Office Salaries $150,000